Dilate the triangle with a scale factor of 2 and the center is (2,3)
Point A (2,4) Point B (-3,3) Point C (2,-2)

Answer by greenestamps(13198)



Point A(2,4) with center (2,3).
Center to point A is 0 units in x direction, 1 unit in y direction.
With scale factor of 2, center to point A' is 0 units in x direction, 2 units in y direction.
A'(2+0,3+2) = (2,5)

Point B(-3,3) with center (2,3).
Center to point B is -5 units in x direction, 0 units in y direction.
With scale factor of 2, center to point B' is -10 units in x direction, 0 units in y direction.
B'(2-10,3+0) = (-8,3)

Point C(2,-2) with center (2,3).
Center to point C is 0 units in x direction, -5 units in y direction.
With scale factor of 2, center to point C' is 0 units in x direction, -10 units in y direction.
C'(2+0,3-10) = (2,-7)